OpenStudy (anonymous):

Solve -10y+18=-3(t-4)=6(t-7)+8 in steps??

OpenStudy (anonymous):

10y+18=3(t-4)=6(t-7)+8 10y=3t-12-18 10y=3t-30 y=(3t-30)/10=(6t-42+8) y = 3t-30 = 60t -340 y = -30+340 = 60t - 3t y=310 = 57t y=310-57t

OpenStudy (anonymous):

Thank you.